If 100 gallons of 75% glycerine solution is made up by combining a 30% glycerine solution with a 90% glycerine solution, how much of each solution must be used?
Answer by Menjax(62)

x=amount of 30% solution
y=amount of 90% solution
x+y=100
100 * 75/100 = x * 30/100 + y * 90/100
simplifying 75=x*3/10+y*9/10
using x=100-y
75=100-y*3/10+y*9/10
75=200-3y/10 + 9y/10
750=200-3y + 9y
550/6=y
550/6 + x = 100
550/6=100-x
550=600-6x
-50=-6x
x=50/6
x=25/3
y=275/3